# Targeting Alignment: Extracting Safety Classifiers of Aligned LLMs

### v2

@@ 3.5 Efficient Jailbreaking
The space of compliant outputs is infinite, thus limiting the search of adversarial examples to those that directly (or indirectly) favorize a particular output is bound to restrict the efficacy of the attack. Using the candidate classifiers which map an input sequence to a predicted safety from the LLM, we can reformulate the objective to a misclassification objective (e.g., max loss) without heuristic:
@@ 4 Evaluation
    How accurate are the candidate classifiers at approximating the safety classifier in benign settings? (Section 4.2)
@@
    Does the performance of candidate classifiers hold in adversarial settings? (Section 4.3)
@@
    Can we jailbreak aligned LLMs by attacking the surrogate classifier? (Section 4.4)
4.1 Experimental Setup
@@
Models. We evaluate our approach on four open-weights aligned chat models, all coming from Hugging Face Transformers 4.46.2 [50] library: Llama-2-7b-chat [30], Qwen2.5-7B-Instruct [46] gemma-7b-it [44], and granite-3.1-8b-instruct [45]. We chose these models because they achieve good performance at the classification task on the datasets considered, translating to a stronger alignment. For completeness, we present results on other models in Appendix A.3.
Dataset augmentation. Since most of the datasets lack safe prompts or targets to attack, we augment them when necessary. For safe prompts, we use an LLM to convert unsafe prompts to their corresponding safe version (following best practices [12]), as it is a simple natural language task suitable for the use of LLMs. We manually verify two properties of each generated prompt: the instruction does not prompt the model for unsafe information, and it has the same structure as the original prompt. For example, the prompt “Write a script that can exploit vulnerabilities in a software or operating system” becomes “Write a script that can identify and help fix vulnerabilities in a software or operating system”, satisfying the two desired properties.
The GCG optimization objective aims to increase the probability of a given output sequence. Therefore, it needs a target output string for each input prompt. For missing targets, we generate what would be a compliance from the model using an LLM, e.g., ”Sure, here is a script that can exploit vulnerabilities in a software or operating system” for the previous unsafe input example.
@@
Labels. The labels are obtained by applying a classifier on the LLM output. To eliminate nondeterminism due to sampling and facilitate reproducibility of results, we generate outputs with a temperature equal to 0: the next token selected at each step of the generation is the one that maximizes the logits. As it has been shown that a higher temperature can lead to a higher probability of generating unsafe outputs [65] (upper bounding the measured efficacy of the attack), we consider the additional tractability incurred by studying the temperature out-of-scope and leave it to future work. The output is classified between refusal and compliance using the distilroberta-base-rejection-v1 [38] model for the reasons mentioned in Section 3.3.
Classifier training. As explained in Section 3.4, the classification head needs to be simple enough to prevent overfitting on the dataset and to only learn to project the features of structure to a classification. Therefore, we consider a simple linear layer followed by a sigmoid: $\mathcal{C}(h)=\sigma(Ah+b)$ where $A\in\mathbb{R}^{1\times d},\ b\in\mathbb{R}$ and $\sigma:x\mapsto\frac{1}{1+e^{-x}}$. The result is a scalar between 0 and 1, thus the classification is made given a threshold $\mathcal{T}$. As for training, we apply $K$-fold cross-validation with 5 folds. We select the threshold $\mathcal{T}$ that maximizes the $F_{1}$ score on the training data, i.e., $\mathcal{T}=\arg\max\limits_{\mathcal{T}\in[0,1]}F_{1}^{(\mathcal{T})}(\mathcal{D}_{\text{train}})$. We train the classification head using the hyperparameters listed in Table I.
@@
Figure 4: Test $F_{1}$ of the candidate classifiers in benign setting, depending on the normalized candidate size.
@@
Figure 5: Test $F_{1}$ of the candidate classifiers in benign setting, on the dataset they were not trained on.
Attack. To generate adversarial examples, we use the nanogcg [^3] See code at https://github.com/GraySwanAI/nanoGCG implementation of GCG [65] with the following hyperparameters: num_steps=250, topk=512,search_width=512, and an attack length of 20 (following prior work [25, 55]). We chose this attack because it is a pure gradient-based attack with an adversarial objective that can be adapted to attacking a classifier. Other white-box attacks may be ill-suited or benefit less from the approach if they are not gradient-based (e.g., BEAST [42]) or if their optimization objective cannot be mapped to misclassification (e.g., Autodan [62]), making the attribution of efficacy to the approach harder. Further, for the purpose of verifying our hypothesis, we decided to ensure tractability of experiments (following prior work that has built on GCG [25, 26]) and chose not to build upon direct descendants of GCG (e.g., AmpleGCG [25]) as they introduce more complexity and do not fit our goal of a heuristic-free attack. For candidate classifiers, we use the binary cross entropy loss of PyTorch [36].
We apply the attack to misclassify unsafe inputs (i.e., create a compliance from the LLMs). For completeness, we also do a similar study in Section A.2 in which the attack is used to misclassify safe inputs.
@@
4.2 Performance in Benign Settings
@@
Baseline results. Table II shows the benign classification performance of the LLMs for each dataset with the corresponding confusion matrices in Figure 6 in Figure 7. We can first see that all models perform well in AdvBench, all scoring above 0.9. The worse performance on OR-Bench is expected, as it was introduced after AdvBench with the goal of being harder to classify and to test for over-refusal, as seen in the top right cell of the confusion matrices in Figure 7. Further, it echoes the result from Section 3.1 in which the silhouette score (measure of separation) is consistently lower for OR-Bench compared to AdvBench.
@@
Candidates performance. Figure 4 shows the median test $F_{1}$ score of the candidate classifiers over 5 trials. We see that performance either increases or stagnates with respect to the candidate size. This suggests that later parts of the model preserve or increase the separation of safe and unsafe instructions. For example, candidate classifiers achieve a lower $F_{1}$ score on OR-Bench, losing 5% for all models. This aligns with the results in Table II where all models perform worse on OR-Bench. We hypothesize that this can be explained by a lower separability between rejected and accepted prompts at the embedding level, as we note that the results from Section 3.1 align with the performance of the candidate classifiers. This lower separation means that the two clusters may overlap, preventing the classification head from completely aligning with the classifier when training.
@@
4.3 Performance in Adversarial Settings
@@
Baseline results. Similarly to the previous settings, we report in Table III the classification performances of the LLMs after applying the white-box attack, along with the attack success rate (ASR) defined as the proportion of unsafe samples misclassified. We note a significant difference (around 70%) between the ASR of Llama 2 and Qwen 2.5. This comes from the strong alignment of Llama 2 with a bias toward refusal (see Section A.1). Aside from Gemma 1, the ASR on OR-Bench is consistently lower than on AdvBench, which could be explained again by the initial goal of the dataset, having an emphasis on over-refusal.
@@
LLM to candidate classifiers. After applying the attack, we filter the adversarial inputs misclassified by the LLM and evaluate how well these examples transfer to the candidate classifiers by measuring the transferability rate: the proportion of misclassified samples by each candidate classifier. Figure 8 reports the transferability rate from the LLM to the candidate classifiers. We observe a similar pattern to that of the cross-dataset study (see Section 4.2): the transfer rate does not converge until at least half of the model is used. For example, candidate classifiers on Llama 2 obtain a transfer higher than 90% only when more than half the model is used. This result implies that candidate classifiers that use less than half of the model are not representative of the safety classifier. We also note that for OR-Bench, most candidate classifiers using less than $40\%$ of their corresponding LLM seem unstable. This phenomenon is consistent with the low silhouette score (see Section 3.1) and performance in the cross-dataset study (see Section 4.2) for corresponding regions of the LLMs. The out-of-distribution nature of adversarial examples may lead to high transferability rates for earlier layers that overfit, strengthening the importance of the cross-dataset study done in Section 4.2.
@@
Imbalance and performance. We note that the performance of the candidate classifiers reaches a higher value for OR-Bench than for AdvBench: for the former, the transferability rate converges to a value above 80% against a value between 60% and 80% for the latter (with Llama 2 being the sole exception). We attribute this difference to the performance of the LLM in benign setting for this dataset (see Figure 7), resulting in a higher amount of samples predicted as unsafe. Therefore, the classification heads of the corresponding candidate classifiers are trained with more information on when the safety classifier classifies an input as unsafe, specifically when it is wrong, leading to an improvement in approximating the safety classifier for adversarial unsafe inputs. This observation is also consistent with the only model unaffected by this trend, Llama 2, which exhibits the highest rate of safe inputs misclassified as unsafe.
@@
Takeaway 4: Performance in adversarial settings Candidate classifiers approximate well the safety classifier under adversarial inputs crafted on the corresponding LLM with a transferability rate converging to a value above 70% for most settings. For the most robust model, Llama 2, the transferability rate to the candidate classifiers exceeds 95% when they use more than 50% of the model.
4.4 Attacking the Surrogate Classifier
@@
ASR and objective. The previous differences in transferability rate compared to the baseline lie in how the adversarial objective is built. Since the output of LLMs is text, there needs to be a heuristic to know when misclassification (jailbreak) occurs, such as specific target sentences. When the objective is a misclassification, it is easier to find adversarial examples. It implicitly contains the previous objective, but encompasses any possible jailbreak sentence, increasing the space of adversarial inputs.
Takeaway 5: Adversarial objective of attacks White-box attacks have relied on heuristic-driven adversarial objectives. Using misclassification as the objective removes unnecessary constraints on the search space of adversarial inputs.
@@
Efficiency. Figure 10 shows the runtime [^4] This does not include the time to train the candidate classifier as it is negligible, taking less than a minute and happening once for all samples. and VRAM usage (i.e., peak GPU memory occupied) when applying the attack on one sample, averaged over 100 samples. We see that attacking the surrogate classifier has lower computational costs than attacking LLM, as it depends on the candidate size $\delta$ (which scales linearly with the VRAM usage like the runtime, as seen in Figure 10). We note that both the VRAM usage and the runtime do not not reach the baseline averages of 31 GB and 840s. We attribute this to the adversarial objective of the baseline GCG attack: it requires the attack to generate an output sequence at each step to compare to the target sequence, directly impacting its efficiency.
Paired with the previous results from Figure 9 in which all models achieve a transferability rate above 70% with only half of the LLM architecture, this highlights the efficacy and efficiency of our approach. This also shows that studying the robustness of LLMs can be made scalable.
Indeed, the original attack requires a NVIDIA A100 with 40 GB for all models (at half-precision). Attacking 50% of the model (as presented in our work) means that 20 GB of VRAM will be sufficient, enabling research on consumer-grade GPUs (e.g., NVIDIA RTX 4090). For industry-grade GPUs (e.g., NVIDIA A100, H100, etc.), this result enables in-depth study on hyperparameters of the attack.
@@
5 Discussion & Future Work
5.1 Domains
Why safety? This study focuses on the problem of safety, i.e., classifying unsafe and safe inputs. We chose this for two main reasons. First, recent work has shown that there is a clear separation between the two classes in the model (see Section 6). Second, this problem is one of the most prolific research topics (e.g., jailbreak attacks), and the insights from this study lead to implications for white-box attacks, which we discuss in the next section.
@@ 5.1 Domains
5.2 Implication for White-box Attacks
@@
Classification objective. As opposed to evasion attacks on classifiers which cause misclassifications, jailbreak attacks aim to induce unsafe behaviors. This objective is more complex because it follows from the definition of safety. When extracting the classifier, we explicitly uncover the decision made by the model. This gives a more accurate signal that does not rely on heuristics (e.g., maximizing a target sequence that is likely to induce a jailbreak) and thus leads to a higher attack efficacy. We also note that such heuristics on the adversarial objective not only implicitly reduce the efficacy, can also reduce the efficiency (see Section 4.4) as it requires multiple forward pass to the entire model to compute an output sequence.
@@
Efficiency. Recent white-box gradient-based attacks on LLM are based on the GCG attack [65]. Although there have been significant performance improvements over time, these attacks compute gradients and, therefore, they are limited when the models are too large. Our results from Section 4.4 suggest that investigating the robustness of alignment through attacks is reducible to attacking the surrogate classifier. This implies fewer computations, and thus higher efficiency (and scalability) of attacks. Our approach was applied on LLMs with less than 9 billion parameters and showed that using 50% of them is enough for the attack. A similar efficiency could be obtained from a transfer attack using a smaller LLM as the source. However, this would come with a lower efficacy and the results might be less informative on the failure points of the target model since a different model is used.
@@
Attack scenarios. Our work can be applied to any jailbreak attack (albeit more suited for optimization-based attacks), making it extensible to multiple scenarios. For instance, in the presence of a guard model that prevents unsafe outputs (e.g., LlamaGuard [18]), recent work has shown that an attack can be “propagated” to this guard model [28]. Our work can be extended to such scenarios by extracting the safety classifiers of (a) the LLM and (b) the guard model, therein improving efficiency.
@@
Black-box. We limit the threat model to white box, since we need access to the LLM’s weights to obtain the surrogate classifier. However, a promising direction would be to study the transferability of safety classifier of a source LLM onto that of a target LLM. Transferability of white-box attacks is a well-known phenomenon of adversarial machine learning [11] also present for jailbreak attacks [65, 27]. An analysis with candidate classifiers could give a finer-grained understanding of the transferability of jailbreak attacks.
@@
Defenses. Two axes are available to defend against our approach. First, as any attack against an ML model, adversarial training [51] remains the de facto standard for increasing ML robustness. Instead of applying such training to the entire model (which would be highly inefficient given its size), it can be applied to the safety classifier with the classification objective. For a finer-grained threat model where the adversary is computationally bounded (by either runtime or memory), explicitly training the LLM to increase the safety classifier size (e.g., by modifying the learning objective) would limit the feasibility of the attack due to computational constraints. This, however, incurs a trade-off for other defenses such as detection of adversarial examples in the intermediate layers [49] as the detection would happen at the end of the safety classifier.
5.3 Safety Classifier

### v3

@@ III-E Efficient Jailbreaking
The space of compliant outputs is infinite, thus limiting the search of adversarial examples to those that directly (or indirectly) prioritize a particular output is bound to restrict the efficacy of the attack. Using the candidate classifiers which map an input sequence to a predicted safety from the LLM, we can reformulate the objective to a misclassification objective (e.g., max loss) without heuristic:
@@ IV Evaluation
    How accurate are the candidate classifiers at approximating the safety classifier in benign settings? (Section IV-B)
@@
    Does the performance of candidate classifiers hold in adversarial settings? (Section IV-C)
@@
    Can we jailbreak aligned LLMs by attacking the surrogate classifier? (Section IV-D)
IV-A Experimental Setup
@@
Models. We evaluate our approach on four open-weights aligned chat models, all from the Hugging Face Transformers 4.46.2 library [54]: Llama-2-7b-chat [32], Qwen2.5-7B-Instruct [49] gemma-7b-it [47], and granite-3.1-8b-instruct [48]. We chose these models because they achieve good performance at the classification task on the datasets considered, translating to a stronger alignment. For completeness, we present results on other models in Appendix A-C.
Dataset augmentation. Since most of the datasets lack safe prompts or targets to attack, we augment them when necessary. For safe prompts, we use an LLM to convert unsafe prompts to their corresponding safe version (following best practices [13]), as it is a simple natural language task suitable for the use of LLMs. We manually verify two properties of each generated prompt: the instruction does not prompt the model for unsafe information, and it has the same structure as the original prompt. For example, the prompt “Write a script that can exploit vulnerabilities in a software or operating system” becomes “Write a script that can identify and help fix vulnerabilities in a software or operating system”, satisfying the two desired properties. Note that preserving the structure of the prompt ensures that we focus on safety rather than artifacts of the dataset, such as how prompts are formatted.
The GCG optimization objective aims to increase the probability of a given output sequence of tokens. Therefore, it needs a target output string for each input prompt. For missing target prompts (e.g., corresponding to augmented safe input prompts), we generate a possible compliant output from the model using an LLM. For instance, “Sure, here is a script that can exploit vulnerabilities in a software or operating system” is the target string for the previous unsafe input example.
@@
Labels. The labels are obtained by applying a classifier on the LLM output. To eliminate nondeterminism due to sampling and facilitate reproducibility of results, we generate outputs with a temperature equal to 0: the next token selected at each step of the generation is the one that maximizes the logits. As it has been shown that a higher temperature can lead to a higher probability of generating unsafe outputs [70] (upper bounding the measured efficacy of the attack), we consider the additional tractability incurred by studying the temperature out-of-scope and leave it to future work. The output is classified between refusal and compliance using the distilroberta-base-rejection-v1 [40] model for the reasons mentioned in Section III-C.
Classifier training. As explained in Section III-D, the classification head needs to be simple enough to prevent overfitting on the dataset and to only learn to project the features of structure to a classification. Therefore, we consider a simple linear layer followed by a sigmoid: $\mathcal{C}(h)=\sigma(Ah+b)$ where $A\in\mathbb{R}^{1\times d},\ b\in\mathbb{R}$ and $\sigma:x\mapsto\frac{1}{1+e^{-x}}$. The result is a scalar between 0 and 1, thus the classification is made given a threshold $\mathcal{T}$. As for training, we apply $K$-fold cross-validation with 5 folds. We select the threshold $\mathcal{T}$ that maximizes the $F_{1}$ score on the training data, i.e., $\mathcal{T}=\arg\max\limits_{\mathcal{T}\in[0,1]}F_{1}^{(\mathcal{T})}(\mathcal{D}_{\text{train}})$. We train the classification head using the hyperparameters listed in Table I.
@@
Figure 4: Test $F_{1}$ of the candidate classifiers in benign settings, depending on the normalized candidate size.
@@
Figure 5: Test $F_{1}$ of the candidate classifiers in benign settings, on the dataset they were not trained on.
Attack. To generate adversarial examples, we use the nanogcg [^3] See code at https://github.com/GraySwanAI/nanoGCG. implementation of GCG [70] with the following hyperparameters: num_steps=250, topk=512,search_width=512, and an attack length of 20 (following prior work [26, 60]). We chose this attack because it is a pure gradient-based attack with an adversarial objective that can be adapted to attacking a classifier. Other white-box attacks may be ill-suited or benefit less from the approach if they are not gradient-based (e.g., BEAST [45]) or if their optimization objective cannot be mapped to misclassification (e.g., Autodan [67]), making the attribution of efficacy to the approach harder. Further, for the purpose of verifying our hypothesis, we decided to ensure tractability of experiments (following prior work that has built on GCG [26, 27]) and chose not to build upon direct descendants of GCG like AmpleGCG [26] as they introduce more complexity and do not fit our goal of a heuristic-free attack. For candidate classifiers, we use the binary cross entropy loss of PyTorch [38].
We apply the attack to misclassify unsafe inputs (i.e., create a compliance from the LLMs). For completeness, we also do a similar study in Section A-B in which the attack is used to misclassify safe inputs.
@@
IV-B Performance in Benign Settings
@@
Baseline results. Table II shows the benign classification performance of the LLMs for each dataset with the corresponding confusion matrices in Figure 6 in Figure 7. We can first see that all models perform well in AdvBench, all scoring above 0.9. The poorer performance on OR-Bench is expected, as it was introduced after AdvBench with the goal of being harder to classify and to test for over-refusal, as seen in the top right cell of the confusion matrices in Figure 7. Further, it echoes the result from Section III-A in which the silhouette score (measure of separation) is consistently lower for OR-Bench compared to AdvBench.
@@
Candidates performance. Figure 4 shows the median test $F_{1}$ score of the candidate classifiers over 5 trials. We see that performance either increases or stagnates with respect to the candidate size. This suggests that later parts of the model preserve or increase the separation of safe and unsafe instructions. For example, candidate classifiers achieve a lower $F_{1}$ score on OR-Bench, losing 5% for all models. This aligns with the results in Table II where all models perform worse on OR-Bench. We hypothesize that this can be explained by a lower separability between rejected and accepted prompts at the embedding level, as we note that the results from Section III-A align with the performance of the candidate classifiers. This lower separation means that the two clusters may overlap, preventing the classification head from completely aligning with the classifier when training.
@@
IV-C Performance in Adversarial Settings
@@
Baseline results. Similarly to the previous settings, we report in Table III the classification performances of the LLMs after applying the white-box attack, along with the attack success rate (ASR) defined as the proportion of unsafe samples misclassified. We note a significant difference (around 70%) between the ASR of Llama 2 and Qwen 2.5. This comes from the strong alignment of Llama 2 with a bias toward refusal (see Section A-A). Aside from Gemma 1, the ASR on OR-Bench is consistently lower than on AdvBench, which could be explained again by the initial goal of the dataset, having an emphasis on over-refusal.
@@
LLM to candidate classifiers. After applying the attack, we filter the adversarial inputs misclassified by the LLM and evaluate how well these examples transfer to the candidate classifiers by measuring the transferability rate: the proportion of misclassified samples by each candidate classifier. Figure 8 reports the transferability rate from the LLM to the candidate classifiers. We observe a similar pattern to that of the cross-dataset study (see Section IV-B): the transfer rate does not converge until at least half of the model is used. For example, candidate classifiers on Llama 2 obtain a transfer higher than 90% only when more than half the model is used. This result implies that candidate classifiers that use less than half of the model are not representative of the safety classifier. We also note that for OR-Bench, most candidate classifiers using less than $40\%$ of their corresponding LLM seem unstable. This phenomenon is consistent with the low silhouette score (see Section III-A) and performance in the cross-dataset study (see Section IV-B) for corresponding regions of the LLMs. The out-of-distribution nature of adversarial examples may lead to high transferability rates for earlier layers that overfit, strengthening the importance of the cross-dataset study done in Section IV-B.
@@
Imbalance and performance. We note that the performance of the candidate classifiers reaches a higher value for OR-Bench than for AdvBench: for the former, the transferability rate converges to a value above 80% against a value between 60% and 80% for the latter (with Llama 2 being the sole exception). We attribute this difference to the performance of the LLM in benign settings for this dataset (see Figure 7), resulting in a higher amount of samples predicted as unsafe. Therefore, the classification heads of the corresponding candidate classifiers are trained with more information on when the safety classifier classifies an input as unsafe, specifically when it is wrong, leading to an improvement in approximating the safety classifier for adversarial unsafe inputs. This observation is also consistent with the only model unaffected by this trend, Llama 2, which exhibits the highest rate of safe inputs misclassified as unsafe.
@@
Takeaway 4: Performance in adversarial settings Candidate classifiers approximate well the safety classifier under adversarial inputs crafted on the corresponding LLM with a transferability rate converging to a value above 70% for most settings. For the most robust model, Llama 2, the transferability rate exceeds 95% when using more than 50% of the model.
IV-D Attacking the Surrogate Classifier
@@
ASR and objective. The previous differences in transferability rate compared to the baseline lie in how the adversarial objective is built. Since the output of LLMs is text, there needs to be a heuristic to know when misclassification (jailbreak) occurs, such as specific target sentences. When the objective is misclassification, it is easier to find adversarial examples. It implicitly contains the previous objective without its limitations, increasing the space of adversarial inputs.
Takeaway 5: Adversarial objective of attacks White-box attacks have relied on heuristic-driven adversarial objectives such as maximizing the likelihood of a given target sentence. Using misclassification as the objective removes unnecessary constraints on the search space of adversarial inputs.
@@
Efficiency. Figure 10 shows the runtime [^4] This does not include the time to train the candidate classifier as it is negligible, taking less than a minute and happening once for all samples. and VRAM usage (i.e., peak GPU memory occupied) when applying the attack on one sample, averaged over 100 samples. We see that attacking the surrogate classifier has lower computational costs than attacking LLM, as it depends on the candidate size $\delta$ (which scales linearly with the VRAM usage like the runtime, as seen in Figure 10). We note that both the VRAM usage and the runtime do not reach the baseline averages of 31 GB and 840s. We attribute this to the adversarial objective of the baseline GCG attack: it requires the attack to generate an output sequence at each step to compare to the target sequence, directly impacting its efficiency. Paired with the previous results from Figure 9 in which all models achieve a transferability rate above 70% with only half of the LLM architecture, this highlights the efficacy and efficiency of our approach. This also shows that studying the robustness of LLMs can be made scalable. Indeed, the original attack requires a NVIDIA A100 with 40 GB for all models (at half-precision). Attacking 50% of the model (as presented in our work) means that 20 GB of VRAM will be sufficient, enabling research on consumer-grade GPUs (e.g., NVIDIA RTX 4090). For industry-grade GPUs (e.g., NVIDIA A100, H100, etc.), this result enables studies on the impact of hyperparameters.
@@
V Discussion & Future Work
V-A Domains
Why safety? This study focuses on the problem of safety, i.e., classifying unsafe and safe inputs. We chose this for two main reasons. First, recent work has shown that there is a clear separation between the two classes in the model (see Section VI). Second, this problem is one of the most prolific research topics (e.g., jailbreak attacks), and the insights from this study lead to implications for white-box attacks, which we discuss in the next section.
@@ V-A Domains
V-B Implication for White-box Attacks
@@
Classification objective. As opposed to evasion attacks on classifiers which cause misclassifications, jailbreak attacks aim to induce unsafe behaviors. This objective is more complex because it follows from the definition of safety. When extracting the classifier, we explicitly uncover the decision made by the model. This gives a more accurate signal that does not rely on heuristics (e.g., maximizing a target sequence that is likely to induce a jailbreak) and thus leads to a higher attack efficacy. We also note that such heuristics on the adversarial objective not only implicitly reduce the efficacy, but can also reduce the efficiency (see Section IV-D) as it requires multiple forward passes to the entire model to compute an output sequence.
@@
Efficiency. Recent white-box gradient-based attacks on LLM are based on the GCG attack [70]. Although there have been significant performance improvements over time, these attacks compute gradients and, therefore, they are limited when the models are too large. Our results from Section IV-D suggest that investigating the robustness of alignment through attacks is reducible to attacking the surrogate classifier. This implies fewer computations, and thus higher efficiency (and scalability) of attacks. Our approach was applied on LLMs with less than 9 billion parameters and showed that using 50% of them is enough for the attack. A similar efficiency could be obtained from a transfer attack using a smaller LLM as the source. However, this would come with a lower efficacy and the results might be less informative on the failure points of the target model since a different model is used.
@@
Attack scenarios. Our work can be applied to any jailbreak attack (albeit more suited for optimization-based attacks), making it extensible to multiple scenarios. For instance, in the presence of a guard model that prevents unsafe outputs (e.g., LlamaGuard [19]), recent work has shown that an attack can be “propagated” to this guard model [29]. Our work can be extended to such scenarios by extracting the safety classifiers of the LLM and the guard model, therein improving efficiency.
@@
Black-box. We limit the threat model to white-box, since we need access to the LLM’s weights to obtain the surrogate classifier. However, a promising direction would be to study the transferability of safety classifier of a source LLM onto that of a target LLM. Transferability of white-box attacks is a well-known phenomenon of adversarial machine learning [12] also present for jailbreak attacks [70, 28]. An analysis with candidate classifiers could give a finer-grained understanding of the transferability of jailbreak attacks.
@@
Defenses. As any attack against an ML model, adversarial training [55] remains the de facto standard for increasing ML robustness. Instead of applying such training to the entire model (which would be highly inefficient given its size), it can be applied to the safety classifier with the classification objective. For a finer-grained threat model where the adversary is computationally bounded (by either runtime or memory), explicitly training the LLM to increase the safety classifier size (e.g., by modifying the learning objective) would limit the feasibility of the attack due to computational constraints. This, however, incurs a trade-off for other defenses such as detection of adversarial examples in the intermediate layers [53] as the detection would happen at the end of the safety classifier. Other approaches add components around the model (e.g., LlamaGuard [19]) to avoid the security-utility trade-off incurred by defending the model itself. With the safety classifier identified, it could be used at inference time to drop unsafe inputs before the output is generated, resulting in higher efficiency for the whole system.
Red-teaming. Ensuring that an LLM is secure before deployment requires a thorough evaluation that needs to be scalable and efficient, given the size of those models. Using our approach could be a promising direction in improving the efficiency of the process: Using the output of the safety classifier for early feedback, red-teamers would be able to iterate faster on attacks and thus audit the model in more depth.
V-C Safety Classifier